# Configurational entropy and stability conditions of fermion and boson stars

### v1

@@ Abstract
In a remarkable study by M. Gleiser and N. Jiang (Phys. Rev. D 92, 044046, 2015), the authors demonstrated that the stability regions of neutron stars, within the framework of the simple Fermi gas model, and self-gravitating configurations of complex scalar field (boson stars) with various self couplings, obtained through traditional perturbation methods, correlates with critical points of the configurational entropy with an accuracy of a few percent. Recently, P. Koliogiannis et al. (Phys. Rev. D 107, 044069 2023) found that while the minimization of the configurational entropy generally anticipates qualitatively the stability point for neutron stars and quark stars, this approach lacks universal validity. In this work, we aim to further elucidate this issue by seeking to reconcile these seemingly contradictory findings. Specifically, we calculate the configurational entropy of bosonic and fermionic systems, described by interacting Fermi and Boson gases, respectively, that form compact objects stabilized by gravity. We investigate whether the minimization of configurational entropy coincides with the stability point of the corresponding compact objects. Our results indicate a strong correlation between the stability points predicted by configurational entropy and those obtained through traditional methods, with the accuracy of this correlation showing a slight dependence on the interaction strength. Consequently, the stability of compact objects, composed of components obeying Fermi or Boson statistics, can alternatively be assessed using the concept of configurational entropy.
@@ I Introduction
In view of the above questions, the main motivation of the present work is to provide an extended examination of the statements of Refs. [23] and [24]. In Ref. [23] the authors found that the stability regions of neutron stars (in the framework of the simple Fermi gas model) as well as of self-gravitating configurations of complex scalar field (boson stars) with various self couplings (detailed reviews in Refs. [42, 43, 44]), obtained from traditional perturbation methods, correlating the critical points of the CE with an accuracy of a few percent. In Ref. [24] the authors found that the suggested prediction of the stability by the minimization of the CE, concerning neutron stars and quark stars, does not have, at least quantitatively, universal validity although in several cases it qualitatively predicts the existence of the stability point. In this work, we aim to further elucidate this issue by seeking to reconcile these seemingly contradictory findings. Specifically, we calculate the configurational entropy of bosonic and fermionic systems, described by interacting Fermi and Boson gases, respectively, that form compact objects stabilized by gravity. We investigate whether the minimization of configurational entropy coincides with the stability point of the corresponding compact objects. Our results indicate a strong correlation between the stability points predicted by configurational entropy and those obtained through traditional methods, with the accuracy of this correlation showing a slight dependence on the interaction strength.
The paper is organized as follows. In Sec. II we present the basic formalism of the hydrodynamic equilibrium and the role of analytical solutions while in Sec. III, we review the definition of the configurational entropy. The parametrization of the equations of state is provided in Sec. IV and in Sec. V the results of the present study are laid out and discussed. Finally, Sec. VI contains the concluding remarks.
@@ II Hydrodynamic equilibrium and analytical solutions
In general, to obtain realistic solutions, it is most natural to numerically solve the TOV equations by incorporating an EoS that describes the relationship between pressure and density within the fluid interior. Alternatively, one can seek analytical solutions to the TOV equations, though these solutions may lack physical relevance. While there are hundreds of analytical solutions to the TOV equations [45, 46], only a few are of significant physical interest. In this work, we employ two of these noteworthy solutions: the Schwarzschild (constant-density interior solution) and the Tolman VII solution [45, 46]. It is important to note that analytical solutions are highly valuable, as they often provide explicit expressions for the quantities of interest and are instrumental in verifying the accuracy of numerical calculations. Below, we briefly describe these two fundamental analytical solutions.
@@
  • Tolman VII solution: The Tolman VII solution has been extensively employed in neutron star studies while its physical realization has been examined, very recently, in detail [47, 48]. The stability of this solution has been examined by Negi et al[49, 50] and also confirmed in Ref. [51]. The energy density and the pressure read as [51]
@@ III Configurational entropy in momentum space
It is notable that the function $F(k)$ in the case of zero momentum, coincides with the gravitational mass of the compact object, that is $F(0)\equiv M$, since by definition (see Eq. (2))
@@
where the density $\rho(r)$ derives from the solution of the TOV equations. Moreover, we define the modal fraction $f({\bf k})$ [23]
@@
and also the function $\tilde{f}({\bf k})=f({\bf k})/f({\bf k})_{\rm max}$, where $f({\bf k})_{\rm max}$ is the maximum fraction, which is given in many cases by the zero mode $k=0$, or by the system’s longest physics mode, $|k_{min}|=\pi/R$. The above normalization guarantees that $\tilde{f}({\bf k})\leq 1$ for all values of ${\bf k}$.
@@ IV Equation of state
The present study focuses on the role of the CE as a stability condition of compact objects obeying Fermi or Boson statistics including a parameterized self-interaction. The compact objects under consideration are neutron stars (introduced by a simplified EoS), boson stars and other type of astrophysical objects composed of fermions or bosons, such as dark matter stars. In following, the equations of interacting Fermi and boson gases are introduced.
@@ IV.1 Interacting Fermi Gas (FG)
where $m_{\chi}$ is the particle mass, which considered equal to $m_{\chi}=939~{}{\rm MeV/c^{2}}$ for reasons of simplicity (this holds throughout the study), $n_{\chi}$ is the number density and
@@ IV.2 Interacting Boson Gas (BG)
To enrich the study, we included the case of compact stars composed of bosonic matter and specifically that of an interacting Boson gas. As the construction of the EoS for the aforementioned gas is not unambiguously defined, we introduce three cases based on different assumptions. It needs to be noted that since the scalar field only vanishes at spatial infinity, boson stars do not have a specific radius where the energy density and pressure vanish. Thus, we do not use a momentum cut-off scheme, $0\leq|{\bf k}|\leq\infty$.
@@
    BG - C1: The first derivation of the EoS of boson stars with a repulsive interaction was given in Ref. [53] and since then it has been used extensively in the corresponding calculations. In particular, the energy density is given as
@@
    where the interaction parameter $z$ is given by $z=u^{2}(\hbar c)^{3}/(m_{\chi}c^{2})^{2}$ (in units of $\rm MeV^{-1}~{}fm^{3}$) and the quantity $u={\rm g}_{\chi}/m_{\phi}c^{2}$ defines the strength of the interaction in analogy to the case of fermions. In this study we have considered the values $z=[0.001,0.005,0.01,0.05,0.1,0.3,0.5]~{}(\rm MeV^{-1}~{}fm^{3})$.
@@
    where the index $n$ is restricted to $n>2$. In this study we have considered the values $n=[4,5]$ for MT (hereafter case (a)) and $n=[3,4,5]$ for VT (hereafter case (b)). An additional reason for using the mentioned cases is because they lead to different mass-radius diagrams (depending on the index values) and thus, cover a large range of cases that may correspond to boson stars.
@@
Figure 1: (a) Gravitational mass as a function of the radius. (b) The corresponding dependence of the gravitational mass as a function of the central density. (c) The corresponding CE as a function of the central density and two analytical solutions of TOV equations (${\rm a}$ and ${\rm b}$ are constants [23]). The inset figure indicates the location of the minimization of the CE. The black diamonds indicate the stability points due to the TM while the open circles correspond to the minimum of the CE. The panels in order correspond to: EoSs of Sec. IV.1, EoSs of Sec. IV.2: case 1, EoSs of Sec. IV.2: case 2, EoSs of Sec. IV.2: case 3.
@@ V Results and Discussion
As a first step, we employed two analytical solutions, namely the Schwarzchild’s and Tolman’s VII solutions, in order to calculate the CE. It is worth pointing out that although it is more natural to use a realistic EoS for the fluid interior in order to solve the Einstein’s field equations, the use of analytical solutions has the advantage that by having an explicit form, the examination of the implied physics becomes simpler. We should remark that the analytical solutions are a source of infinite number of EoSs (plausible or not). Consequently, they can be used extensively, to introduce and establish some universal approximations (for more details and discussion see also Ref. [51]). In both cases, we formulate the dependence in the form
@@
In the case of the FG, the first panel of Fig. 1 displays that the points due to TM and CE are located in close proximity, validating the CE method for the location of the stability point. However, a detailed presentation of the percentage error on some interesting quantities, namely the gravitational mass, the radius, the central energy density and the compactness, as shown in Table 1, signal a different behavior. While the error in the gravitational mass is lower than 5%, the error in the central energy density can reach up to almost 99%, depending each time to the value of the interaction. The latter have its origin in the creation of a plateau immediately after the rapid decrease of the CE. The existence of a plateau maintains the CE in a narrow region, while the central energy density is spanning in a wide region. In addition, as the data indicate, there is no simple relation between the interaction and the corresponding error to establish a pattern. Thus, in the FG case, while the CE can potentially establish the maximum gravitational mass with good accuracy, the proper description of the central energy density is almost impossible. It needs to be noted that for some specific values of the interaction, the location of the total minimum in CE was not successful, even at high values of densities beyond the maximum mass configuration (unstable region). In that cases, we located a local minimum near the density that corresponds to the maximum mass configuration. The aforementioned statement holds for all cases under consideration in the present study.
@@
In the case of the BG, a similar behavior with the FG is observed. Once more, Fig. 1 in the second, third and fourth panel, displays a visually small difference between the TM and CE points. Nevertheless, Table 1 illustrates that the error in the underlying quantities aligns with the trend observed in the FG. Specifically, the error in gravitational mass is less than 5%, while the error in central energy density can extend up to 100%. In this instance as well, the CE can be employed to determine the maximum mass, but not the associated central energy density. The observed behavior in both the FG and BG cases strengthens the argument that the location of the stability point is an intrinsic property of the EoS.
@@
As one might easily suspect the CE is related to the star’s compactness, Fig. 2 displays the dependence of the CE on the compactness for (a) FG, (b) BG - C1, (c) BG - C2, and (d) BG - C3. In the compactness plane the CE creates also a plateau, similar to the central energy density plane in Fig .1(c), but in the majority of the cases it is not so extended, depending each time on the corresponding mass-radius diagram of Fig.1(a). In the FG case, the error can reach values up to 12%, where in the majority of the cases the error is lower than 6%. This result is in accordance with the error in the gravitational mass along with the corresponding error values in the radius. As the radius is more sensitive to the structure of the star, this sensitivity is also presented in the error, reaching values close to 15%. Concerning the BG case, the error in the compactness is established in general under 12% and in the corresponding radius under 14%.
@@
For a visual presentation of Table 1 for the predictions of the two methods, Fig. 3 displays the percentage error on the predictions for various fermionic and bosonic stars properties (gravitational mass, radius, central density and compactness) as a function of the relative parameterization of the interaction strength $(y,w,z,n)$ in each particular case. As a general comment, the convergence of the two methods is clearly better in the case of fermion stars compared to that of boson stars (at least for the selected range of parameterizations). Furthermore, for the case of a boson star, the choice of the EoS is decisive for the accuracy of the convergence of the two methods.
@@ VI Concluding Remarks
Compact objects composed of either fermionic or bosonic matter have been employed for studying the configurational entropy as a mean of stability. The aforementioned quantity should be in alignment with the stability criterion of TM method. It is important to note that the existence of a stable configuration is a property of gravity and independent of the EoS. However, the specific location of the stability point is influenced by the underlying EoS [57]. Considering the above points, one might expect, as suggested in Ref. [23], that the minimization of the CE is a consequence of gravity within the framework of general relativity. If this is the case, the relevant minimum (which should be a total minimum) should be independent of the applied EoS. However, our findings indicate that this is not universally true for compact objects.
The CE was studied in light of the gravitational mass, radius, central energy density and compactness. The TM and the CE method for the location of the stability point converge, with a good or moderate accuracy, for the three out of four quantities under consideration. In fact, the most accurate prediction lies with the maximum mass, where the difference reaches values lower than 5%. In addition, a quantity with good accuracy is also the corresponding radius with errors up to 15%, while in the majority of the cases, the error is lower than 10%. As a result, the combination of the aforementioned macroscopic quantities, which is the compactness, is also a quantity with accuracy lying on values lower than 12%. These three quantities are decisive indicators of the macroscopic quantities of compact stars and leading to the ultimate result that in a macroscopic scale, the two methods for locating the stability point, are in agreement.
The last quantity under consideration, which is the central energy density, leads to enormous amounts of error that can reach values up to 100%. From this point of view, the two methods contradict each other rendering the central energy density unreliably calculated through the CE method.
Figure 2: Configurational entropy as a function of the compactness parameter. The black diamonds indicate the stability points due to the TM while the open circles correspond to the minimum of the CE. The panels in order correspond to: (a) EoSs of Sec. IV.1, (b) EoSs of Sec. IV.2: case 1, (c) EoSs of Sec. IV.2: case 2, (d) EoSs of Sec. IV.2: case 3.
Figure 3: Percentage error in gravitational mass (diamonds), radius (squares), central energy density (plus signs) and compactness (triangles) as a function of the interaction strength parameter of the CE method with respect to the TM method. The panels in order correspond to: (a) EoSs of Sec. IV.1, (b) EoSs of Sec. IV.2: case 1, (c) EoSs of Sec. IV.2: case 2, (d) EoSs of Sec. IV.2: case 3.

### v2

@@ Abstract
In a remarkable study by M. Gleiser and N. Jiang [Phys. Rev. D 92, 044046, 2015], the authors demonstrated that the stability regions of neutron stars, within the framework of the simple Fermi gas model, and self-gravitating configurations of complex scalar field (boson stars) with various self couplings, obtained through traditional perturbation methods, correlate with critical points of the configurational entropy with an accuracy of a few percent. Recently, P. Koliogiannis et al. [Phys. Rev. D 107, 044069 2023] found that while the minimization of the configurational entropy generally anticipates qualitatively the stability point for neutron stars and quark stars, this approach lacks universal validity. In this work, we aim to further elucidate this issue by seeking to reconcile these seemingly contradictory findings. Specifically, we calculate the configurational entropy of bosonic and fermionic systems, described by interacting Fermi and boson gases, respectively, that form compact objects stabilized by gravity. We investigate whether the minimization of configurational entropy coincides with the stability point of the corresponding compact objects. Our results indicate a strong correlation between the stability points predicted by configurational entropy and those obtained through traditional methods, with the accuracy of this correlation showing a slight dependence on the interaction strength. Consequently, the stability of compact objects, composed of components obeying Fermi or boson statistics, can alternatively be assessed using the concept of configurational entropy.
@@ I Introduction
In view of the above questions, the main motivation of the present work is to provide an extended examination of the statements of Refs. [23] and [24]. In Ref. [23] the authors found that the stability regions of neutron stars (in the framework of the simple Fermi gas model) as well as of self-gravitating configurations of complex scalar field (boson stars) with various self couplings (detailed reviews in Refs. [42, 43, 44]), obtained from traditional perturbation methods, correlate the critical points of the CE with an accuracy of a few percent. In Ref. [24] the authors found that the suggested prediction of the stability by the minimization of the CE, concerning neutron stars and quark stars, does not have, at least quantitatively, universal validity although in several cases it qualitatively predicts the existence of the stability point. In this work, we aim to further elucidate this issue by seeking to reconcile these seemingly contradictory findings. Specifically, we calculate the configurational entropy of bosonic and fermionic systems, described by interacting Fermi and boson gases, respectively, that form compact objects stabilized by gravity. We investigate whether the minimization of configurational entropy coincides with the stability point of the corresponding compact objects. Our results indicate a strong correlation between the stability points predicted by configurational entropy and those obtained through traditional methods, with the accuracy of this correlation showing a slight dependence on the interaction strength.
The paper is organized as follows. In Sec. II we present the basic formalism of the hydrodynamic equilibrium and the role of analytical solutions while in Sec. III, we review the definition of the configurational entropy. The parametrization of the equations of state is provided in Sec. IV and and the results of the present study are laid out and discussed in Sec. V. Finally, Sec. VI contains the concluding remarks.
@@ II Hydrodynamic equilibrium and analytical solutions
In general, to obtain realistic solutions, it is most natural to numerically solve the TOV equations by incorporating an EoS that describes the relationship between pressure and energy density within the fluid interior. Alternatively, one can seek analytical solutions to the TOV equations, though these solutions may lack physical relevance. While there are hundreds of analytical solutions to the TOV equations [45, 46], only a few are of significant physical interest. In this work, we employ two of these noteworthy solutions: the Schwarzschild (constant-density interior solution) and the Tolman-VII solution [45, 46]. It is important to note that analytical solutions are highly valuable, as they often provide explicit expressions for the quantities of interest and are instrumental in verifying the accuracy of numerical calculations. Below, we briefly describe these two fundamental analytical solutions.
@@
  • Tolman-VII solution: The Tolman-VII solution has been extensively employed in neutron star studies while its physical realization has been examined, very recently, in detail [47, 48]. The stability of this solution has been examined by Negi et al[49, 50] and also confirmed in Ref. [51]. The energy density and the pressure read as [51]
@@ III Configurational entropy in momentum space
It is notable that the function $F({\bf k})$ in the case of zero momentum, coincides with the gravitational mass of the compact object, that is $F(0)\equiv M$, since by definition (see Eq. (2))
@@
Moreover, we define the modal fraction $f({\bf k})$ [23]
@@
and also the function $\tilde{f}({\bf k})=f({\bf k})/f({\bf k})_{\rm max}$, where $f({\bf k})_{\rm max}$ is the maximum fraction, which is given in many cases by the zero mode $k=0$, or by the system’s longest physics mode, $|k_{\rm min}|=\pi/R$. The above normalization guarantees that $\tilde{f}({\bf k})\leq 1$ for all values of ${\bf k}$.
@@ IV Equation of state
The present study focuses on the role of the CE as a stability condition of compact objects obeying Fermi or boson statistics including a parametrized self-interaction. The compact objects under consideration are neutron stars (introduced by a simplified EoS), boson stars and other type of astrophysical objects composed of fermions or bosons, such as dark matter stars. In following, the equations of interacting Fermi and boson gases are introduced.
@@ IV.1 Interacting Fermi gas (FG)
where $m_{\chi}$ is the particle mass, considered equal to $m_{\chi}=939.565~{}{\rm MeV/c^{2}}$ for reasons of simplicity (this holds throughout the study), $n_{\chi}$ is the number density and
@@ IV.2 Interacting boson gas (BG)
To enrich our study, we included the case of compact stars composed of bosonic matter and specifically that of an interacting boson gas. As the construction of the EoS for the aforementioned gas is not unambiguously defined, we introduce three cases based on different assumptions. It needs to be noted that since the scalar field only vanishes at spatial infinity, boson stars do not have a specific radius where the energy density and pressure vanish. Thus, we do not use a momentum cut-off scheme, $0\leq|{\bf k}|\leq\infty$.
@@
    BG-C1: The first derivation of the EoS of boson stars with a repulsive interaction was given in Refs. [53, 54, 55] and since then it has been used extensively in the corresponding calculations. In particular, the energy density is given as
@@
    with the interaction parameter $z$ being equal to $z=u^{2}(\hbar c)^{3}/(m_{\chi}c^{2})^{2}$ (in units of $\rm MeV^{-1}~{}fm^{3}$) and the quantity $u={\rm g}_{\chi}/m_{\phi}c^{2}$ defines the strength of the interaction in analogy to the case of fermions. In this study we have considered the values $z=[0.001,0.005,0.01,0.05,0.1,0.3,0.5]~{}(\rm MeV^{-1}~{}fm^{3})$.
@@
    where the index $n$ is restricted to $n>2$. In this study we have considered the values $n=[4,5]$ for MT [hereafter case (a)] and $n=[3,4,5]$ for VT [hereafter case (b)]. An additional reason for using the mentioned cases is because they lead to different mass-radius diagrams (depending on the index values) and thus, cover a large range of cases that may correspond to boson stars.
@@
Figure 1: (a) Gravitational mass as a function of the radius. (b) The corresponding dependence of the gravitational mass as a function of the central density. (c) The corresponding CE as a function of the central density and two analytical solutions of TOV equations, where their constants $\rm a$ and $\rm b$ are given in the text below Eq. (16) and ensure the dimensionless units of the two axes. The black diamonds indicate the stability points due to the TM while the open circles correspond to the minimum of the CE. The panels in order correspond to FG EoSs, BG-C1 EoSs, BG-C2 EoSs, BG-C3 EoSs.
@@ V Results and Discussion
As a first step, we employed two analytical solutions, namely the Schwarzchild’s and Tolman’s-VII solutions, in order to calculate the CE. It is worth pointing out that although it is more natural to use a realistic EoS for the fluid interior in order to solve the Einstein’s field equations, the use of analytical solutions has the advantage that by having an explicit form, the examination of the implied physics becomes simpler. We should remark that the analytical solutions are a source of infinite number of EoSs (plausible or not). Consequently, they can be used extensively, to introduce and establish some universal approximations (for more details and discussion see also Ref. [51]). In both cases, we formulate the dependence in the form
@@
In the case of the FG, the first panel of Fig. 1 displays that the points due to TM and CE are located in close proximity, validating the CE method for the location of the stability point. However, a detailed presentation of the percentage error on some interesting quantities, namely the gravitational mass, the radius, the central density and the compactness, as shown in Table 1, signal a different behavior. While the error in the gravitational mass is lower than 5%, the error in the central density can reach up to almost 99%, depending each time to the value of the interaction. The latter have its origin in the creation of a plateau immediately after the rapid decrease of the CE. The existence of a plateau maintains the CE in a narrow region, while the central density is spanning in a wide region. In addition, as the data indicate, there is no simple relation between the interaction and the corresponding error to establish a pattern. Thus, in the FG case, while the CE can potentially establish the maximum gravitational mass with good accuracy, the proper description of the central density is almost impossible. It needs to be noted that for some specific values of the interaction, the location of the total minimum in CE was not successful, even at high values of densities beyond the maximum mass configuration (unstable region). In that cases, we located a local minimum near the density that corresponds to the maximum mass configuration. The aforementioned statement holds for all cases under consideration in the present study.
@@
In the case of the BG, a similar behavior with the FG is observed. Once more, Fig. 1 in the second, third and fourth panel, displays a visually small difference between the TM and CE points. Nevertheless, Table 1 illustrates that the error in the underlying quantities aligns with the trend observed in the FG. Specifically, the error in gravitational mass is less than 5%, while the error in central density can extend up to 100%. In this instance as well, the CE can be employed to determine the maximum mass, but not the associated central density. The observed behavior in both the FG and BG cases strengthens the argument that the location of the stability point is an intrinsic property of the EoS.
@@
As anticipated, the CE is related to the star’s compactness as illustrated in Fig. 2, which displays the dependence of the CE on the compactness for (a) FG, (b) BG-C1, (c) BG-C2, and (d) BG-C3. In the compactness plane the CE creates also a plateau, similar to the central density plane in Fig .1(c), but in the majority of the cases it is not so extended, depending each time on the corresponding mass-radius diagram of Fig.1(a). In the FG case, the error can reach values up to 12%, where in the majority of the cases the error is lower than 6%. This result is in accordance with the error in the gravitational mass along with the corresponding error values in the radius. As the radius is more sensitive to the structure of the star, this sensitivity is also presented in the error, reaching values close to 15%. Concerning the BG case, the error in the compactness is established in general under 12% and in the corresponding radius under 14%.
Figure 2: Configurational entropy as a function of the compactness parameter. The black diamonds indicate the stability points due to the TM while the open circles correspond to the minimum of the CE. The panels in order correspond to (a) FG EoSs, (b) BG-C1 EoSs, (c) BG-C2 EoSs, (d) BG-C3 EoSs.
Figure 3: Percentage error in gravitational mass (diamonds), radius (squares), central density (plus signs) and compactness (triangles) as a function of the interaction strength parameter of the CE method with respect to the TM method. The panels in order correspond to: (a) FG EoSs, (b) BG-C1 EoSs, (c) BG-C2 EoSs, (d) BG-C3 EoSs.
@@
For a visual presentation of Table 1 for the predictions of the two methods, Fig. 3 displays the percentage error on the predictions for various fermionic and bosonic stars properties (gravitational mass, radius, central density and compactness) as a function of the relative parametrization of the interaction strength $(y,w,z,n)$ in each particular case. As a general comment, the convergence of the two methods is clearly better in the case of fermion stars compared to that of boson stars (at least for the selected range of parametrizations). Furthermore, for the case of a boson star, the choice of the EoS is decisive for the accuracy of the convergence of the two methods.
As it is evident, one of the primary distinctions between fermionic and bosonic systems lies in the choice of the momentum cut-off scheme. In this context, we present a detailed analysis of its role for the systems under consideration. Specifically, we examine two cut-off schemes: (a) $k_{\rm min}=\pi/R$, which is based on the radius of the boson star originated from the condition that the pressure vanishes at the star’s surface, $P(R)=0$, and (b) $k_{\rm min}=\pi/R_{\rm eff}$, which is the effective radius where most of the star’s mass is concentrated and defined as [23]
[…]
The two schemes, along with the initial case where $k_{\rm min}=0$, are compared in detail in Fig. 4, for the BG-C1 (0.001) EoS as a representative case. The introduction of a cut-off scheme significantly improves the accuracy of the results, with case (a) showing noteworthy improvement in accuracy over the no cut-off case. Furthermore, the cut-off scheme in case (b) alters the behavior of the CE, leading to a sharper and more distinct outcome. The comparative results from the different cut-off schemes are also summarized in Table 2, demonstrating that the choice of the momentum cut-off contributes significantly to the convergence of the two methods under consideration.
Figure 4: (a) Gravitational mass as a function of the radius for the BG-C1 (0.001) EoS. (b) The corresponding dependence of the gravitational mass as a function of the central density. (c) The corresponding CE as a function of the central density and two analytical solutions of TOV equations, where their constants $\rm a$ and $\rm b$ are given in the text below Eq. (16) and ensure the dimensionless units of the two axes. The black diamonds indicate the stability points due to the TM while the open markers correspond to the minimum of the CE with: (a) no cut-off ($k_{\rm min}=0$; circle), (b) cut-off at $k_{\rm min}=\pi/R$ (cross), and (c) cut-off at $k_{\rm min}=\pi/R_{\rm eff}$ (square).
In conclusion, beyond the importance of the EoS, the selection of the appropriate cut-off scheme in the momentum plays a critical role in improving the accuracy and convergence of the two methods.
[…]
$\rm w$ $\rm M$ $\rm R$ $\rm\rho_{c}$ $\rm\beta$
BG-C1 No cut-off 0.001 1.776 9.012 58.881 7.953
$k_{\rm min}=\pi/R$ 0.001 0.068 1.762 9.083 1.724
$k_{\rm min}=\pi/R_{\rm eff}$ 0.001 1.140 7.158 28.926 7.744
@@ VI Concluding Remarks
Compact objects composed of either fermionic or bosonic matter have been employed for studying the configurational entropy as a means of stability. The aforementioned quantity should be in alignment with the stability criterion of TM method. It is important to note that the existence of a stable configuration is a property of gravity and independent of the EoS. However, the specific location of the stability point is influenced by the underlying EoS [59]. Considering the above points, one might expect, as suggested in Ref. [23], that the minimization of the CE is a consequence of gravity within the framework of general relativity. If this is the case, the relevant minimum (which should be a total minimum) should be independent of the applied EoS. However, our findings indicate that this is not universally true for compact objects.
The CE was studied in light of the gravitational mass, radius, central density and compactness. The TM and the CE method for the location of the stability point converge, with a good or moderate accuracy, for the three out of four quantities under consideration. In fact, the most accurate prediction lies with the maximum mass, where the difference reaches values lower than 5%. In addition, a quantity with good accuracy is also the corresponding radius with errors up to 15%, while in the majority of the cases, the error is lower than 10%. As a result, the combination of the aforementioned macroscopic quantities, which is the compactness, is also a quantity with accuracy lying on values lower than 12%. These three quantities are decisive indicators of the macroscopic quantities of compact stars and leading to the ultimate result that in a macroscopic scale, the two methods for locating the stability point, are in agreement.
The last quantity under consideration, which is the central density, leads to enormous amounts of error that can reach values up to 100%. From this point of view, the two methods contradict each other rendering the central density unreliably calculated through the CE method. However, it is important to note that, while the previous analysis may result in significant errors in certain cases, the situation is different for boson stars. By carefully selecting an appropriate momentum cut-off, the convergence between the two methods can be greatly improved across all four quantities under consideration. This highlights the critical role of the momentum cut-off in mitigating errors and enhancing the overall accuracy of the results.